# DSNS: The Deep Space Network Simulator

### v1

@@ I-A Requirements
xxeoverse [14] Surrey Python Unreleased LEO megaconstellation simulator based on Mininet [15]. Real-time simulation of large networks including dynamic ISLs.
@@
SatScope [17] NUDT China Python Unreleased 2025 LEO constellation network simulator based on VTK with a focus on satellite internet routing/coverage.
@@ I-B Contributions
In this paper we present the Deep Space Network Simulator (DSNS): a new network simulator optimized for interplanetary networks that satisfies all the requirements established above. In contrast to existing offerings, explored in Section II, DSNS scales well to large numbers of nodes, supports arbitrary network topologies with interplanetary links, and uses an underlying event-based simulation to enable simulations to run faster than real time. Furthermore, DSNS is easily extensible thanks to its modular architecture – new protocols can be added, removed, and swapped out, and new layers of the network stack can be implemented by simply defining new sets of rules upon which messages and events are matched.

### v3

@@ I-A Requirements
xxeoverse [14] Surrey Python Non-commercial 2025 LEO megaconstellation simulator based on Mininet [15]. Real-time simulation of large networks including dynamic ISLs.
@@
SatScope [17] NUDT China Python Non-commercial 2025 LEO constellation network simulator based on VTK with a focus on satellite internet routing/coverage.
@@ I-B Contributions
In this paper we present the Deep Space Network Simulator (DSNS): a new network simulator optimized for interplanetary networks that satisfies all the requirements established above. In contrast to existing offerings, explored in Section II, DSNS scales well to large numbers of nodes, supports arbitrary network topologies with interplanetary links, and uses an underlying event-based simulation to enable simulations to run faster than real time. Furthermore, DSNS is easily extensible thanks to its modular architecture – new protocols can be added, removed, and swapped out, and new layers of the network stack can be implemented by simply defining new sets of rules upon which messages and events are matched. To facilitate future research, DSNS has been made fully open source, released under the GNU GPLv3 license. [^1] The source code and documentation can be found at https://github.com/ssloxford/DSNS.
@@ VI-D Physical Layer
VI-E Security
Finally, we discuss the ways in which DSNS can be used to enable security research in this area. In the source code for the simulator we provide mechanisms for simulating the following attacks:
  • Global message loss: We provide configurable randomized message loss across the whole network through the LossConfig class, enabling the testing of reliable delivery and rerouting mechanisms under realistic loss conditions.
  • Targeted message loss: The LossConfig also supports per-link loss rates, enabling simulation of more targeted attacks and ensuring redelivery and rerouting still works under direct attack.
  • Link flooding attacks: Finally, we provide tools to simulate targeted flooding attacks, denying service by exhausting the available bandwidth on a link. This can be accessed via the TrafficFloodActor, and can be configured to start and end at a specific time, target a single link or a group of links, and use different amounts of bandwidth.
Furthermore, the simplicity and extensibility of DSNS makes it easy to build further attack mechanisms, making it particularly useful to assess the behavior of different protocols and strategies under a wide range of attacks.